Suppose that the cost of producing n pairs of shoes is given by the expression C(n)=30+5n. The average cost of producing a pair of shoes is given by A(n)=C(n)/n. What is the average cost at n=30?
Answer #1

Average cost = A(n)=C(n)/n = (30+5n)/n

So, when n= 30, Average cost = (30+5n)/n = (30 + 5 *30)/30 = 180/30 = 6.

So, when n= 30, average cost = 6.